### v2

# Johan Håstad

Abstract
We study Frege proofs for the functional and onto graph Pigeon Hole Principle defined on the $n\times n$ grid where $n$ is odd. We are interested in the case where each formula in the proof is a depth $d$ formula in the basis given by $\land$, $\lor$, and $\neg$. We prove that in this situation the proof needs to be of size exponential in $n^{\Omega(1/d)}$. If we restrict the size of each line in the proof to be of size $M$ then the number of lines needed is exponential in $n/(\log M)^{O(d)}$. The main technical component of the proofs is to design a new family of random restrictions and to prove the appropriate switching lemmas.
1 Introduction
In this paper we study formal proofs of formulas in Boolean variables encoding natural combinatorial principles. We can think of these as tautologies but it is often more convenient to think of them as contradictions. When a certain formula, $F$, is a contradiction then its negation, $\bar{F}$ is a tautology and we do not distinguish the two. In particular in the below discussion we might call something a tautology that the reader, possibly rightly, thinks of as a contradiction. We are equally liberal with usage of the word “proof” which might more accurately be called “a derivation of contradiction”.
We are given a set of local constraints that we call “axioms”. These are locally satisfiable but not globally in that there is no global assignment that satisfies all the axioms. A proof derives consequences of the axioms and it is complete when it reaches an obvious contradiction such as 1=0 or that an empty clause contains a true literal.
[…]
A fundamental and popular case is resolution corresponding to $d=1$, where each formula is a disjunction of literals. It is far from easy to analyze resolution but this proof system has been studied for a long time and many questions are now resolved. We do not want to discuss the history of resolution but as it is very relevant for the current paper let us mention that an early milestone was obtained by Haken [7] in 1985 when he proved that the Pigeon Hole Principle (PHP) requires exponential size resolution proofs. The PHP states that $n+1$ pigeons can fly to $n$ holes such that no two pigeons fly to the same hole. There are many ways to code this statement using Boolean variables and we use the basic variant which has $(n+1)n$ Boolean variables $x_{ij}$. Such a variable is true iff pigeon $i$ flies to the hole $j$. The axioms say that for each $i$ there is a value of $j$ such that $x_{ij}$ is true and for each $j$ there is at most one $i$ such that $x_{ij}$ is true. This is clearly a contradiction but to prove this counting is useful and resolution is not very efficient when it comes to counting.
Previous work. The focus of this paper is the more powerful proof system obtained for larger values of $d$ and here a pioneering result was obtained by Ajtai [1] proving superpolynomial lower bounds for the size of any proof for PHP for any fixed constant $d$. The lower bounds of Ajtai were not explicit and Bellantoni et al [3] gave the first such bounds, namely that depth $\Omega(\log^{*}n)$ is needed for the size of the proof to be polynomial. This was greatly improved in two independent works by Krajíček, Pudlák, and Woods [13] and Pitassi, Beame, and Impagliazzo [16], respectively. These two papers established lower bounds for the size of any proof of the PHP of the form exponential in $n^{c^{-d}}$ where $c>1$, and gave non-trivial bounds for depths as high as $\Theta(\log{\log n})$.
Related questions were studied in circuit complexity where the central question is to study the size of a circuit needed to compute a particular function. Here a sequence of results [6],[20], [23], [8] established size lower bounds of the form exponential in $n^{\Omega(1/d)}$ and obtained strong lower bounds for $d$ as large as $\Theta(\log n/\log{\log n})$. The results were obtained for the parity function and here it is easy to show that this function can be computed by circuits of matching size. To see that PHP allows proof of size exponential in $n^{O(1/d)}$ is more difficult but was established in 2001 by Atserias et al [2].
A technique used in many of these papers is called “restrictions”. The idea is simply to, in a more or less clever way, give values to most of the variables in the object under study and to analyze the effect. One must preserve [^1] One does not really preserve a function or a formula and an object of size $n$ is reduced to a similar object of size $f(n)$ for some $f(n)<n$. the function computed (or tautology being proved) while at the same time be able to simplify the circuits assumed to compute the function or the formulas in the claimed proof. An important reason that the lower bounds in circuit complexity were stronger than those in proof complexity is that it is easier to preserve a single function than an entire tautology with many axioms. Due to these complications strong lower bounds for depths beyond $\Theta(\log{\log n})$, for any tautology, remained unknown for several decades.
The first result that broke this barrier was obtained by Pitassi et al [17] who obtained super-polynomial lower bounds for depths up to any $o(\sqrt{\log n})$. The tautology considered was first studied by Tseitin [21] and considers a set of linear equations modulo two defined by a graph. The underlying graph for [17] is an expander. These results were later extended to depth $\Theta(\log n/\log{\log n})$ by Håstad [10] and in this case the underlying graph is the two-dimensional square grid. The bounds obtained were further improved by Håstad and Risse [11].
All results mentioned so far only discuss total size. For resolution, each formula derived is a clause and hence of size at most $n$ but for other proof systems it is interesting to study the number of lines in the proof and the sizes of lines separately. Pitassi, Ramakrishnan and Tan [15] had the great insight that a technical strengthening of the used methods yields much stronger bounds for this measure than implied by the size bounds. They combined some of the techniques of [10] with some methods from [17] to establish that if each line is of size at most $M$ then the number of lines in a proof that establishes the Tseitin principle over the square grid needs to be exponential in $n2^{-d\sqrt{\log M}}$. By using some additional ideas Håstad and Risse [11] fully extended the techniques of [10] to this setting improving the bounds to exponential in $n/(\log M)^{O(d)}$.
Our results. Despite this progress, the thirty year old question whether the PHP allows polynomial size proof of depth $O(\log\log n)$ remained open. The purpose of this paper is to prove that it does not and that lower bounds similar to those for the Tseitin tautology also apply to the PHP. To build on previous techniques we study what is known as the graph PHP where the underlying graph is an odd size two-dimensional grid.
As the side length of the grid is odd, if one colors it as a chess board, the corners are of the same color and let us assume this is white. In the graph PHP on the grid, there is a pigeon on each white square and it should fly to one of the adjacent black squares that define the holes. This graph PHP is the result of the general PHP where most variables are forced to take the value 0. Each pigeon is only given at most four alternatives. Clearly any proof for the general PHP can be modified to give a proof of the graph PHP by replacing some variables by the constant false. To limit ourselves further we prove our lower bounds for what is known as the functional and onto PHP. We add axioms saying that each pigeon can only fly to one hole and each hole receives exactly one pigeon.
Phrased slightly differently, the functional and onto PHP on the grid says that there is a perfect matching of the odd size grid and we heavily use local matchings. We can compare this to the Tseitin tautology on the grid studied by [10, 15, 11] that states that it is possible to assign Boolean values to the edges of the grid such that there is an odd number of true variables next to any node. As a perfect matching would immediately yield such an assignment, the PHP is a stronger statement and possibly easier to refute. In particular, any lower bound for functional and onto PHP gives the same lower bound for the Tseitin contradiction. We do not, however, think of the results of the current paper as a new proofs for the results in [11] (with slightly weaker bounds). In fact, as the construction of this paper shares many features with the construction of [11], it is better to think the current proof as an adjustment of the latter proof taking care of complications that we here only work with matchings. Let us turn to discuss the main technical point, namely to prove a “switching lemma”.
By assigning values to most variables in a formula it is possible to switch a small depth-two formula from being a CNF to being a DNF and the other way around. In the basic switching lemma used to prove circuit lower bounds [6, 23, 8], uniformly random constant values are substituted for a majority of the variables. Such restrictions are the easiest to analyze but are less useful in proof complexity as they do not preserve any interesting tautology.
To preserve a tautology or a complicated function it is useful to replace several old variables by the same new variable, possibly negated. It is possible to be even more liberal and allow old variables to be replaced by slightly more complicated expressions in the new variable. This technique was first introduced explicitly by Rossman, Servedio, and Tan [19] when studying the depth hierarchy for small-depth circuits but had been used in more primitive form in earlier papers. We use such generalized restrictions in this paper.
Multi-switching. The technical strengthening needed by [15] that we discussed above is to improve the standard switching lemma to what is commonly known as a multi-switching lemma. This concept was first introduced independently by Håstad [9] and Impagliazzo, Mathews, and Paturi [12] to study the correlation of small-depths circuits and simple functions such as parity.
In this setting one considers many formulas $(F^{i})_{i=1}^{m}$ and the goal is to switch them all simultaneously in the following sense. There is a small depth (common) decision tree such that at any leaf of the tree it is possible to represent each $F^{i}$ by a small formula of the other type. It was the insight that multi-switching could be used in the proof complexity setting that made it possible for [15] to derive the strong bounds on the number of lines in a proof when each line is short. The extension to multi-switching turns out to be mostly technical and not require any really new ideas. In view of this let us for the rest of this introduction only discuss standard switching.
Techniques used. The most novel part is to design this new space of restrictions that preserve the functional and onto PHP. It has many similarities with the space introduced in [10] and from a very high level point of view, the proofs follow the same path. At the more detailed level in this paper we work with partial matchings of the grid which is a more rigid object than assignment that only satisfy the Tseitin condition of an odd number of true variables next to any node. This results in considerable changes in the details and as a result we get slightly worse bounds.
Once the space of restrictions is in place, two tasks remain. Namely to prove the switching lemmas and then use these bounds to derive the claimed bounds on proof size. This latter part hardly changes compared to previous papers. The switching lemmas follow the same pattern as in [10] and [11] but we need some new combinatorial lemmas and the fact that we work with matchings calls for some modifications.
Outline of the paper. We start with some preliminaries and recall some facts from previous papers in Section 2. We introduce our new space of random restrictions in Section 3. The basic switching lemma is proved in Section 4 and we use it to establish the lower bound for proof size in Section 5. We give the multi-switch lemma in Section 6 and use it, in Section 7, to derive the lower bounds on the number of lines in a proof. We end with some very brief comments in Section 8.
2 Preliminaries
In this section we give some basic definitions and derive some simple properties. We also recall some useful facts from related papers.
2.1 The formula to refute
We study the functional and onto PHP on the odd size $n\times n$ grid. Nodes are given indexed by $(i,j)$ where $1\leq i,j\leq n$ and a node is connected to other nodes where one of the two coordinates is the same and the other differ by 1. As opposed to some previous papers studying the Tseitin tautology, it is here important that we are on the grid and not on the torus as we want a bipartite graph. For any edge, $e=(v,w)$ of adjacent nodes, $v$ and $w$, we have a variable $x_{e}$. The axioms say that for each $v$ if you consider the, at most four, variables adjacent to $v$ exactly one of them is true. This is easy to write as a 4-CNF. The axioms look exactly the same for holes and pigeons.
This formula is a stronger statement compared to the standard PHP formula with $(n^{2}+1)/2$ pigeons and $(n^{2}-1)/2$ holes. Our parameters are slightly different from the standard parameters where $n$ is the number of holes, but we trust the reader to keep this in mind.
We assume that the grid is colored as a chess board and that the corners are white. Thus pigeons correspond to white nodes and holes to black nodes. We prefer to use the colors to distinguish the two as we later use larger objects that inherit the same colors.
2.2 Frege proofs
[…]
  • (Excluded middle) $(p\lor\neg p)$
  • (Expansion rule) $(p\rightarrow p\lor q)$
  • (Contraction rule) $(p\lor p)\rightarrow p$
  • (Association rule) $p\lor(q\lor r)\rightarrow(p\lor q)\lor r$
  • (Cut rule) $p\lor q,\neg p\lor r\rightarrow q\lor r$.
2.3 Decision trees and partial matchings of the grid
The use of decision trees is central to the current paper let us define some notions. In a standard decision tree, at each node we ask for the value of an input variable. In this paper we instead allow queries of the form.
“To which node is $v$ matched?”
The answer to this question determines the value of any variable next to $v$ and thus is more powerful than a single variable question. On the other hand it can be simulated by asking ordinary variable questions for three variables around $v$. Thus within a factor of three in the number of questions, this type of questions is equivalent to variable queries.
We require that the answers along any branch in a decision are locally consistent and keep all such branches. The key property is that it is possible to extend the partial matching given by the answers on this branch to a large fraction of the grid. Product sets of collections of disjoint sets of intervals are convenient for us. The size of an interval is the number of elements it contains. The size of a partial matching is the number of edges it contains.
Definition 2.1
A partial matching $M$ on the $n\times n$ grid of size $t$ is locally consistent if it can be extended to a complete matching of a larger set $S\times T$. We require that each of $S$ and $T$ is the union of disjoint even-sized intervals such that the total size of all intervals is at most $48t$.
Remark. The constant $48$ in the definition is not optimal and with a more careful analysis it can be decreased, but we have opted for shorter proofs rather than optimal constants.
We use two properties from locally consistent matchings. Firstly that given any node not matched in the matching then it is possible to include also this node with a suitable partner. This is proved in Lemma 2.2 below. The second property we need is that a sub-matching of any locally consistent matching is also locally consistent. This is proved in Lemma 2.3. We do not exclude that there is a simpler definition of locally consistent that achieve these two properties, but our definition is one possibility.
Lemma 2.2
Suppose we a locally consistent partial matching, $M$ of size at most $n/50-9$ in the $n\times n$ grid and that we are given a node $v$ not matched by $M$. It is then possible to find a partner, $w$, of $v$, such that $M$ jointly with $(u,w)$ is a locally consistent matching.
Proof:  If $v$ is already in $S\times T$ we can use the same extension. Suppose $v=(a,b)$ where $a\in S$ and $b\not\in T$. It is easy to find $b^{\prime}$ such that $T\cup\{b,b^{\prime}\}$ is a union of even size intervals. Now we can add matchings of $S\times b$ and $S\times b^{\prime}$ using that $S$ is a union of even size intervals.
The case when $a\not\in S$ and $b\in T$ is symmetric and let us handle the case $a\not\in S$ and $b\not\in T$. We can find $b^{\prime}$ as in the previous case, enlarging $T$ to $T^{\prime}=T\cup\{b,b^{\prime}\}$ and then proceed by adding $a$ and a suitable $a^{\prime}$ to $S$.
We enlarge the sizes of $S$ and $T$ by at most 2 but since the size of the matching increases by one this is not a problem.
Let us next establish natural property that any sub-matching of a locally consistent matching is locally consistent. With the given definition of locally consistent, the proof needs some work.
Lemma 2.3
Suppose $M$ is a locally consistent matching. Then any subset of $M$ is locally consistent.
Proof:  Suppose the matching $M$ is of size $t$ and let $M_{0}$ be a sub-matching of size $t_{0}$. We have sets of $S$ of $T$ and an extension of $M$ to a matching of $S\times T$. This, of course, is also an extension of $M_{0}$ and if both these sets are of size at most $48t_{0}$ we are done, so assume that this is not the case and we need to decrease the sizes of $S$ and $T$.
We need to take a closer look on which sub-areas of the grid allow matchings. This is traditionally stated as “tiling an area in the plane with dominos of size 2”. We need to recall some facts about such tilings and we use results from [14] and [5]. For the convenience of the reader we give the argument from scratch.
A “figure” is a number of black and white squares. An example is given in Figure 1. Each square has four edges along its sides and two squares are connected if they share an edge (and not if they only meet at a corner). Connected squares are thus of different colors. We are interested in connected figures but not necessarily simply connected. Each figure has a perimeter which is the set of edges which only belong to one square of the figure. There is an outside perimeter which is a simple path and if the figure is not simply connected, we have perimeter(s) around any interior hole(s).
Figure 1: A figure that cannot be tiled.
It is convenient to use directed paths and we direct the perimeter of a figure in the clockwise order and the perimeters of the holes in the counter-clockwise order. In each case, the figure lies to the right of each perimeter edge when the edge is traversed in its intended direction. For notational convenience we let the coordinates of the corners of the squares be the integer points.
Definition 2.4
The cost of a directed edge is 1 if the square to its right (when traversed in the intended direction) is white and -1 otherwise. The cost of a path is the sum of the costs of its edge. We denote the cost of a path $P$ by $c(P)$.
The following lemma is one reason why the given definition is useful.
Lemma 2.5
Assume we direct the perimeter as described, i.e. the outer perimeter in the clockwise direction and the perimeter around any holes in the counter-clockwise direction. If the figure contains $b$ black squares and $w$ white squares then the total cost of the perimeter is $4(w-b)$.
Proof:  For each square in the figure consider the four edges of the perimeter of this square when it is traversed in the clockwise order. The total cost of all edges is then $4(w-b)$ as each white square contributes 4 and each black square -4. The edges internal to the figure are traversed once in each direction and these two contributions cancel each other. The edges traversed once give the cost of the perimeter and hence the lemma follows.
Suppose we are given a figure, $A$, with equally many black and white squares and we want to determine whether it admits a matching. As we can treat each connected component separately we can assume that $A$ is connected. We use Hall’s marriage theorem to study this question. Let $L$ be a set of black squares and $N(L)$ we the set of white neighbors, both within $A$. It follows from Hall’s marriage theorem that a necessary and sufficient condition for $A$ to be matchable is that for any $L$ we have $|N(L)|\geq L$. For a set $L$ that potentially violate this condition let us look at the figure $A_{0}=L\cup N(L)$. This contains more black squares than white squares and hence the cost of its perimeter must be negative.
This perimeter contains edges from the perimeter of $A$ and some edges internal to $A$. Note that the latter edges all have cost one as there must be a white square to its right. Indeed if there was a black square to its right, $N(L)$ would contain also the square to its left and hence the edge would not belong to the perimeter. This property implies that these edges internal to $A$ must alternate in the sense that every other edges is vertical and every other edge is horizontal. We call such a sub-figure “white bounded”. Using this terminology we have the following lemma which is a combination of the above reasoning and Lemma 2.5.
Lemma 2.6
A figure with an equal number of black and white squares is matchable unless it contains a connected white-bounded sub-figure with a perimeter of negative cost.
For instance in Figure 1 one such sub-figure is bounded by the three white square going down the diagonal in the middle. This sub-figure contains 6 black and 5 white squares.
Let us return to our matching $M_{0}$ which can be extended to a matching of $S\times T$ which are of total size $48t$. The aim is to decrease this size. In order to do this we define a new notion. Let $I=[a,b]$ be an interval and $K$ a multiset contained in the interior of $I$. Define the “left function” of $I$ as follows.
  • $f^{I}_{l}(a)=0$.
  • For $a<i\leq b$ set $f^{I}_{l}(i)=f^{I}_{l}(i-1)+\delta_{i}$ where $\delta_{i}=2$ if $i\not\in K$ and otherwise $\delta_{i}=-k_{i}/2$ where $k_{i}$ is the number of copies of $i$ in $K$.
Similarly we have a right function, defined by $f^{I}_{r}(b)=0$ and $f^{I}_{r}(i)=f^{I}_{r}(i+1)+\delta_{i}$.
Definition 2.7
A set of intervals, $S$, well covers a multi-set $K$ if $S$ covers $K$ and for each $I\in S$ we have $f_{l}^{I}(i)\geq 0$ and $f_{r}^{I}(i)\geq 0$ for each $i\in I$.
We have the following lemma.
Lemma 2.8
Any set $K$ of cardinality $t$ can be well covered by a collection of intervals of even length of total size $6t$.
Proof:  We add the points of $K$ one by one and build the set of intervals at the same time maintaining the property that the current intervals form a well cover of the so far added points. For each point in $K$ we add at most 5 points to $S$. We make the intervals of even length in the end.
We start with $K$ empty and no intervals. Suppose we want to add $p$ to $K$. If $p$ is not in $S$, we add it. We find the two points to the left of $p$ which are as close as possible to $p$ and which are currently not in $S$ and add them to $S$. Similarly we add two points to the right of $p$.
We claim that the resulting set well covers $K$ and establish this by induction. Clearly this is true if $K$ is a singleton and we need to study adding a general point $p$ to $K$.
As intervals not touched by the change remain the same, and left and right are symmetric, it is sufficient to look at $f^{I}_{l}$ for the interval, $I$, that the added point $p$ ends up in. The interval $I$ is the union of some intervals $I^{i}$ existing before the addition of $p$, jointly with $p$ and the four added points.
Define a function $f$ on $I$ by setting $f(x)=f_{l}^{I^{i}}(x)$ if $x\in I^{i}$ and $f(x)=0$ otherwise. We claim that $f_{l}^{I}(x)\geq f(x)+\delta_{x}$ where, for $x$ strictly to the left of $p$, $\delta_{x}$ is twice the number of points added strictly to the left $x$. If $x$ equals $p$ or is to the right of $p$ then $\delta_{x}=0$. Clearly this claim is sufficient to establish that $f_{l}^{I}$ remains non-negative.
The claim uses the alternative characterization of $f^{I}_{l}(x)$ as being twice the number of to left of $x$ that belongs to $I\setminus K$ minus half the number of points in $I\cap K$ to the left of or equal to $x$. If there was no merger of interval the bound on $f_{l}^{I}$ is obvious. That mergers do not decrease $f_{l}^{I}$ follows from the fact that the contribution to $f_{l}^{I}$ from $I^{i}$ is non-negative due to the non-negativity of $f_{l}^{I^{i}}$. We conclude that the claim holds.
Finally, we add a point to any interval in $S$ of odd size. As each interval is of length at least 5, this multiplies its length by at most $6/5$.
The squares of $M_{0}$ gives a number of connected figures and consider any point on the perimeter of one of these figures. We define $K_{1}$ to be the set of all first coordinates of these points, considered as a multiset. This is clearly a set of size at most $8t_{0}$. We let $S_{0}$ be the set given by Lemma 2.8 that well covers $K_{1}$. Similarly considering the second coordinates we get a set $K_{2}$ and let $T_{0}$ be a well cover of this set. If either $S_{0}$ or $T_{0}$ contains points outside $[n]$ we drop these point(s). By Lemma 2.8 both $S_{0}$ and $T_{0}$ are of size at most $48t_{0}$ and hence Lemma 2.3 follows from Lemma 2.9 below.
Lemma 2.9
$M_{0}$ can be extended to a matching of $S_{0}\times T_{0}$.
Proof:  Define the figure $A$ to be $S_{0}\times T_{0}$ with the squares of $M_{0}$ removed. Suppose for contradiction that $A$ cannot be tiled with dominos of size 2. Then by Lemma 2.6 we get a white-bounded connected sub-figure, $A_{0}$, with a perimeter of negative cost. As we only analyze $A_{0}$, we only need to consider its connected component in $S_{0}\times T_{0}$ and hence we can assume that $S_{0}$ and $T_{0}$ both are given by single intervals. The perimeter of $A_{0}$ contains an outer boundary and the perimeters of its holes (if any). First we observe that cost of the perimeter of any hole is non-negative. Suppose on the contrary that you have a hole $H$ with negative cost perimeter.
To have any chance to have a negative cost the boundary $H$ hole must contain some elements of $M_{0}$ and these give a matching $M_{0}^{\prime}$. Now consider any rectangle, $R$ with even side lengths that strictly contains $H$. We claim that that there is no extension of $M_{0}^{\prime}$ to cover all of $R$. This follows as the figure obtained by removing from $H$ from $R$ has a negative cost boundary. Furthermore, this figure is of the form $L\cup N(L)$ for a suitable $L$ and its existence would contradict that $M_{0}^{\prime}$ can be extended to any rectangle with even side lengths that contains its component in $S\times T$.
This proves that no hole in $A_{0}$ contributes a negative cost and next we establish that also the boundary perimeter of $A_{0}$ is of non-negative cost.
If this boundary contains no part of the boundary of $S_{0}\times T_{0}$ then we can apply the above argument and conclude that the cost is positive. On the other hand if it is identical to the boundary of $S_{0}\times T_{0}$ the cost 0.
In the remaining case we can divide this perimeter into segments $s_{i}$ numbered form 00 to $12t-$ 1 where even numbered segments are along the perimeter of $S_{0}\times T_{0}$ and odd segments contain edges that either are in the interior of $A_{0}$ or along the perimeter of $M_{0}$. First note that $c(s_{2j})\geq 0$ for any $j$. This follows as they start and end with a edge of cost 1. Costs of edges along the perimeter alternate between 1 and -1 except at the corners where we get two adjacent edges of the same cost.
Now let us look at the odd numbered components.
Lemma 2.10
For each odd $i$ we have $c(s_{i})\geq 0$.
Proof:  Suppose that $s_{i}$ starts at point $(a,b)$. Then either $a$ is in the interior of $S_{0}$ or $b$ is in the interior of $T_{0}$ and as the two cases are symmetric we may assume the former. Let us for the time being assume that $b$ is the smallest element in $T_{0}$. If $s_{i}$ does not contain any edges from the perimeter of a component of $M_{0}$ each edge on the segment has cost 1 and the lemma is true. Otherwise let $p$ be the last point of $s_{i}$ that is on the perimeter of $M_{0}$. We claim that the cost of the part of $s_{i}$ from $(a,b)$ to $p$ is non-negative. This establishes the lemma as all edges after $p$ on $s_{i}$ are in the interior of $A_{0}$ and hence of cost 1. Let $b^{\prime}$ be the highest value of any point on the segment from $(a,b)$ to $p$. The following claim, of which we postpone the proof, finishes the proof, in the case when $b$ is the smallest element of $T_{0}$.
Claim 2.11
The path from $(a,b)$ to $p$ has cost at least $f_{l}^{T_{0}}(b^{\prime})$.
The case when $b$ is the largest element in $T_{0}$ is completely analogous. We define $b^{\prime}$ to be the lowest value of the second component of the segment and use $f_{r}^{T_{0}}$ in the corresponding claim. This completes the proof of Lemma 2.10 modulo the claim.
Let us establish Claim 2.11. Look at the projection of the given path on the second coordinate. Consider any $b^{\prime\prime}$ with $b<b^{\prime\prime}\leq b^{\prime}$ such that no point of $M_{0}$ has second coordinate $b^{\prime\prime}$. We want to associate cost two with any such value. As no point in $M_{0}$ projects on $b^{\prime\prime}$ any edge with a second coordinate $b^{\prime\prime}$ must be in the interior of $A_{0}$. As previously observed in this part we alternate between vertical and horizontal edges.
Thus there exists one edge with both end points with second coordinate equal to $b^{\prime\prime}$. We associate the cost 1 to $b^{\prime\prime}$ due to this edge. We must also have two edges with different second coordinates and one end point with second coordinate $b^{\prime\prime}$. In most situation there is one with second coordinate $b^{\prime\prime}-1$ and one with second coordinate $b^{\prime\prime}+1$. The only exception is when $b^{\prime\prime}=b^{\prime}$ and this points does not belong to $M_{0}$ in which case both other and points have second coordinate $b^{\prime\prime}-1$. For each of these two edges we associate cost 1/2 with $b^{\prime\prime}$ for a total cost of two associated with $b^{\prime\prime}$. Note also that the total cost caused by any edge in the interior is 1.
The only edges on the path from $(a,b)$ to $p$ that give value -1 are those from the perimeter of $M_{0}$. We associate the cost -1/2 with each of the second coordinates of its end-points. These associated costs follow closely the definition of $f_{l}^{T_{0}}$ and the claim follows by simple inspection. Having established the claim we conclude that the perimeter of the figure $A_{0}$ cannot have negative costs. This contradiction completes the proof of Lemma 2.9.
We are also interesting in matching areas of the plane which are almost a complete squares. The difference is that some squares next to the perimeter are removed. We call such a removed square a “dent”.
Figure 2: A square with 3 white and 3 black dents´.
Lemma 2.12
Suppose have square with even side length and which has the same number of white and black dents. Suppose further that there are at most $M$ dents and no dent is within $M$ of a corner. In this situation there is always a matching of the square. If we have a square with odd side length with one more white node, then the similar statement is true assuming we have one more white dent.
Proof:  We could use the general condition for the existence of a matching given in Lemma 2.6 but as the situation is very structured we find a direct proof to be easier to follow. It also gives better constants (even though, of course, this is not very important). Note first that as the side lenght is even and we have equally many white and black dents, the number of white squares in the figure equals the number of black squares.
Let us first do the case of even side length and start by scetching the argument. We construct a matching in an onion-like fashion using as many pairs along the perimeter as possible. If two dents are at an even distance then it is possible to perfectly match all nodes between the two dents. This happens iff two adjacent dents are of different colors. If two adjacent dents are of the same color we need to match one node with the node towards the interior creating a new dent in the new square which has side lenght two less than the original square. Let us say this more formally.
Suppose the side length is $S$. Take any dent, and assume for concreteness that it is white. Start matching nodes two and two along the perimeter starting with the node next to this dent. This is straightforward until we hit the next dent. If this dent is black we get a perfect match along the perimeter while if it is white we are forced to create a new white dent. We continue on the other side of this dent and go all around the square. We get a new square of side length $S-2$ and some dents. A dent remains iff it is the same color as the dent preceding it.
As we have both black and white dents, the number of dents has decreased by at least two and the distance to the corner has decreased by at most 2. As each matched pair contains one black square and one white square the number of white dents equals the number of black dents. We repeat this process until there are no remaining dents. The rest is easy to match.
If the side length is odd then the process stops when there is only one white dent on the boundary. Also in this case the remaining square is easy to match.
It is easy to see that in most situation where we apply Lemma 2.12 there are many possible matchings. It will be convenient for us to think of the matching as given by the boundary conditions. This can be done in many ways and any deterministic procedure to go from the set of dents to the matching works for us.
Lemma 2.2 ensures that in any decision tree with question of the form “To which node is $v$ matched?” there is one possible answer such that the partial matching created by the path so far results in a locally consistent matching. By simply erasing any branch which is not locally consistent we maintain the property of each branch being locally consistent.
Lemma 2.3 ensures a slightly more subtle property. Suppose we are given a decision tree, $T$, of small depth and we have a small partial matching, $\tau$ and we want to explore $T$ given $\tau$. Lemma 2.2 makes sure that as long as the sum of the size of $\tau$ and the depth of the decision tree is small, it is always possible to find a locally consistent partner for node queried by $T$. Lemma 2.3 makes sure that any such answer is also present in $T$ and hence it has not already been pruned away. Let us state this as a lemma
Lemma 2.13
Let $T$ be a decision tree of depth $t$ and $\tau$ a locally consistent matching of size $t^{\prime}$. Then, provided $t+t^{\prime}\leq n/50$ the decision tree obtained from pruning all branches of $T$ which are not consistent with $\tau$ is a non-empty decision tree.
We use this lemma in many places without explicitly referring to it. We denote the resulting decision tree by $T\lceil_{\tau}$.
We say that a decision tree is a 1-tree if all its leafs are labeled one and similarly we have 0-trees. It might seem redundant to allow such trees but when doing operations on decision trees they naturally occur.
2.4 $t$-evaluations
The concept of $t$-evaluations was introduced by Krajíček et al. [13] and is a very convenient tool for proving lower bounds on proof size. Here we follow the presentation of Urquhart and Fu [22] while using the notation of [10] and [11]. A $t$-evaluation, $\varphi$, is a map from a set of formulas to decision trees of depth at most $t$. We always have the property that this set of formulas is closed under taking sub-formulas. We want the following properties.
  1. 1.
    The constant true is represented by a 1-tree and the constant 0 is represented by a 0-tree.
  2. 2.
    If $F$ is an axiom of the PHP contradiction then $\varphi(F)$ is a 1-tree.
  3. 3.
    If $F$ is a single variable then $\varphi(F)$ is the natural decision tree of depth 1 defining the value of this variable.
  4. 4.
    If $\varphi(F)=T$ then $\varphi(\neg F)$ is a decision tree with the same internal structure as $T$ but where the value at each leaf is negated.
  5. 5.
    Suppose $F=\lor F_{i}$. Consider a leaf in $\varphi(F)$ and the assignment, $\tau$ leading to this leaf. If the leaf is labeled 0 then for each $i$ $\varphi(F_{i})\lceil_{\tau}$ is a 0-tree and if the leaf is labeled 1 then for some $i$, $\varphi(F_{i})\lceil_{\tau}$ is a 1-tree.
The key property for $t$-evaluations is the following lemma.
Lemma 2.14
Suppose we have a derivation using the rules of Section 2.2 starting with the axioms of the functional onto PHP on the $n\times n$ grid. Let $\Gamma$ be the set of all sub-formulas of this derivation and suppose there is a $t$-evaluation whose range includes $\Gamma$ where $t\leq n/150$. Then each line in the derivation is mapped to a 1-tree. In particular we do not reach a contradiction.
Proof:  This is the lemma that corresponds to Lemma 6.4 of [10]. It relies on two properties, namely that each axiom is represented by a 1-tree, and that the derivation rules preserve this property. The first property is ensured by the local consistency of any branch in a decision tree. The second property follows from the fact that the derivation rules are sound and we never “get stuck” in a decision tree. By this we mean that it always possible to continue a branch in a decision tree keeping the values locally consistent. This is ensured by Lemma 2.2 and Lemma 2.3.
The proof is by induction over the number of lines in the proof and we need to discuss each of the five rules. As this is rather tedious we here only discuss the Cut rule which is the most interesting rule. We are confident that the reader can handle the other rules and if this is not the case, the other cases are discussed in detail in the proof of Lemma 6.4 of [10].
The application of the cut rule derives $F=q\lor r$ from $p\lor q$ and $\neg p\lor r$. By induction $\varphi(p\lor q)$ and $\varphi(\neg p\lor r)$ are both 1-trees and we must establish that so is $\varphi(F)$. For contradiction, take a supposed leaf with label 0 in $\varphi(F)$ and let $\tau$ be the assignment leading to this leaf. We know that $\varphi(q)\lceil_{\tau}$ and $\varphi(r)\lceil_{\tau}$ are both 0-trees. Consider any path in $\varphi(p)\lceil_{\tau}$ and let $\tau_{1}$ be the assignment of this path. Assume this leaf is labeled 0, the other case being similar. Now take any path in $\varphi(p\lor q)\lceil_{\tau\tau_{1}}$. As this is a 1-tree the label at this path must be 1. This contradicts that $\varphi(p)\lceil_{\tau_{1}}$ as well as $\varphi(q)\lceil_{\tau}$ are both 0-trees.
We need sets of three assignments of size $t$ to extendable in a locally consistent way and $493t\leq n/$ 49, ensures this is possible.
When studying the number of lines in a proof where each line is short, an extension of Lemma 2.14 is needed. This was first done in [15] and we rely on the argument in the full version of [11].
In this situation it is not possible to have a single $t$-evaluation whose range is all the sub-formulas that appear in the proof. Instead we will have separate $t$-evaluations for each line. Furthermore, for each line these $t$-evaluations live at the leaves of a decision tree. As these are very similar to what is called $\ell$-common decision trees (as defined in [9]) we call them $\ell$-common $t$-evaluations.
Definition 2.15
A set of formulas $(F_{i})_{i=1}^{M}$ has an $\ell$-common $t$-evaluation if there is a decision tree of depth $\ell$ with the following properties. Take any leaf of this decision tree and the $\tau$ be the partial assignment defined by this path. At this leaf we have a $t$-evaluation $\varphi_{\tau}$ of the formulas $(F^{i}\lceil_{\tau}))_{i=1}^{M}$.
We need an extension of Lemma 2.14. Suppose we are given a proof of PHP and for each line, $\lambda$, let $\Gamma_{\lambda}$ be the set of all sub-formulas of the formula appearing in this line and let $\varphi_{\tau}^{\lambda}$ be the $t$-evaluation given at the leaf defined by $\tau$ that contains $\Gamma_{\lambda}$ it is domain.
Lemma 2.16
Suppose we have a derivation using the rules of Section 2.2 starting with the axioms of the functional onto PHP on the $n\times n$ grid. Suppose each $\Gamma_{\lambda}$ allows a $\ell$-common $t$-evaluation. Suppose $\ell+t\leq n/150$, then each line of the derivation is mapped to a 1-tree in all leaves of its decision tree.
The proof of this not very complicated but needs some notation. The important property is that the copies of the same formula appearing in multiple lines are mapped in a consistent way. Take a formula $F$ appearing on lines $\lambda_{1}$ and $\lambda_{2}$. Take any leaf in the common decision tree of $\lambda_{1}$ defined by an assignment $\tau_{1}$. Take any assignment $\tau_{1}^{\prime}$ consistent with $\tau_{1}$ in the decision tree $\varphi_{\tau_{1}}^{\lambda_{1}}(F)$ leading to a branch labeled $b$. We claim the following.
Lemma 2.17
For any $F$, $\tau_{1}$, $\tau_{1}^{\prime}$ as described above consider the decision tree $T=\varphi_{\tau_{2}}^{\lambda_{2}}(F)$ where $\tau_{2}$ is locally consistent with $\tau_{1}$ and $\tau_{1}^{\prime}$. Then any path in $T$ defined by an assignment $\tau_{2}^{\prime}$ consistent with $\tau_{1}$, $\tau_{1}^{\prime}$ and $\tau_{2}$ leads to a leaf labeled $b$.
Proof:  We prove this by induction over the complexity of $F$. The lemma is true when $F$ is a single variable as this follows from the definition of a $t$-evaluation.
Now for the induction step. When $F$ is of the form $\neg F^{\prime}$, then the the lemma follows from the basic properties of $t$-evaluations and the inductive case that the lemma is true for $F^{\prime}$.
Finally assume that $F=\lor_{i=1}^{m}F_{i}$ and suppose it violates the lemma getting a tuple $\tau_{1},\tau_{1}^{\prime},\tau_{2}$, and $\tau_{2}^{\prime}$ leading to leaves with different values in the two trees. Assume with without loss of generality that $b=1$, the situation being symmetric between the two lines.
By the the definition of $t$-evaluations there must be an $i$ such that $\varphi_{\tau_{1}}^{\lambda_{1}}(F_{i})\lceil_{\tau_{1}^{\prime}}$ is a 1-tree. Similarly for $\lambda_{2}$, we know that $\varphi_{\tau_{2}}^{\lambda_{2}}(F_{i})\lceil_{\tau_{2}^{\prime}}$ is a 0-tree. As the four-tuple of restrictions is consistent and $F_{i}$ is simpler than $F$ this contradicts the inductive assumption.
Here we need two sets of pairs of assignments to be locally consistent and hence $t+\ell\leq n/100$ is sufficient.
Lemma 2.17 tells us that the same formula when considered on different lines and under different $t$-evaluations behaves the same. The proof of Lemma 2.16 is now essentially identical to the proof of Lemma 2.14. In this situation we need triplets of three pairs of assignments to be consistent and hence we use $t+\ell\leq n/150$. We omit the details.
3 Restrictions
As stated in the introduction we use a slightly more complicated object than a restriction which normally only gives values to some variables. A restriction in our setting fixes many variables to constants but also substitutes the same variable or its negation for some variables. In a few cases an old variable is substituted by a small logical formula which is a disjunction of size at most three.
We are given an instance of the PHP on the $n\times n$ grid and a restriction, for a suitable parameter $T$, reduces it to a smaller instance on the $(n/T)\times(n/T)$ grid where $(n/T)$ is an odd number.
We divide the grid in to $(n/T)^{2}$ squares, called super-squares, each with side length $T$. Inside each super-square there are $\Delta$ (for a parameter to be fixed) smaller squares that we from now on call “mini-squares”. We pick one mini-square inside each super-square and let these represent the smaller instance. Each super-square has a color as given by a chess board coloring of the reduced instance. For instance the corner super-squares are all white. Each mini-square has the color of its super-square.
Between each mini-square, $s_{i}$ and any mini-square $s_{j}^{\prime}$ in an adjacent super-square we have $3R$ (for a parameter to be chosen) edge disjoint paths, each of even length. We can match exactly all vertices on this path by matching each node of the path to the appropriate adjacent node. We are also interested in matching each node to the other neighbor on the path and in case we need to include one node in each of the two mini-squares to which it is attached creating a dent in the sense of Lemma 2.12. We think of this as using the path as an augmenting path and hence we sometimes refer to these paths as “augmenting paths”. A matching on such a path is of type 0 if it does not include any node from the attached mini-squares and otherwise it is of type 1.
We group the $3R$ paths in groups of three and within each group, two attach at a node which is the same color as the color as the mini-square, while the third one attaches to a node of opposite color. As each path is of even length, the nodes of attachment at the two end-points are of different colors but so are the mini-squares to which the path attaches. Hence the attachment points are either both the same color as the respective mini-square or both the opposite color. The first two augmenting paths may be of type 0 or 1. The third path is always of type 1 and as these paths play little role in the argument we mostly ignore them from now on.
For each augmenting path $P$ we have a corresponding Boolean variable $y_{P}$ which indicates whether it is of type 0 or type 1. By the $R$ fixed paths of type 1 we can conclude that for any mini-square, if exactly half of its varying paths are of type 0 (and hence the other half is type 1) then its interior has equally many white nodes as black nodes remaining.
We set up our restrictions such that it is uniquely determined by the values for the variables $y_{P}$. Outside the paths and and the mini-squares we more or less have a fixed matching. This concludes the high level description of a restriction and let us repeat the argument, now giving all formal details.
3.1 Details of mini-squares and paths
It is convenient to use the concept of a brick which is a square of size $30\times 30$. We think of routing paths through bricks, but in concrete terms such a path is given by two points of attachment of different colors. It the path is of type 0 these two potential dents are matched inside the brick and if the path is of type 1 these two points are matched to the appropriate neighbor in the neighboring brick. We make sure that the there are at most three dents in each side of the brick and of distance at least 12 to any corner and hence we can use Lemma 2.12 to find a matching of the interior. This matching might not look like a path going through the brick but from an abstract point this is a good way to see it.
We have two parameters, $R$ and $\Delta$ and we think of them as follows. $R$ is needed to ensure that of $R$ (almost) random bits it is likely that about half of them are true. For this reason $R=\Theta(\log n)$. The parameter $\Delta$ is picked to make sure that $150R\Delta^{2}$ is slightly less than $T$.
All but one mini-square have side length $120\Delta R$. To the exterior we have $4\Delta R$ bricks on each side and the interesting part is the middle $2\Delta R$ bricks. Half of these are used to route the $3\Delta R$ paths to the $\Delta$ mini-squares in the super-square in the given direction. The other $\Delta R$ middle bricks are used to make sure that there is enough space to route the paths through different blocks between the mini-squares. The bricks closer to the corners are only used to get the distance to the corners needed by Lemma 2.12.
We have a special single mini-square in the top left corner super-square [^2] This could be any white super-square and choosing this particular super-square is just to make some fixed choice.. It has side length $1120\Delta R+$ 1 which in particular is an odd number but is otherwise like the other mini-squares. We call this the “designated survivor”.
$T$$T$
Figure 3: The placement of mini-squares and super-squares and some paths. The designated survivor is the checkered mini-square. The matchings along top row and leftmost column are indicated by solid lines.
In the top row, the $n-120R\Delta-1$ nodes outside the designated survivor are matched in a horizontal matching. Similarly the nodes in the leftmost column outside the designated survivor are matched in a vertical matching. This essentially eliminates one row and one column outside the top left super-square and we assume for convenience that $n\equiv 1$ modulo $30$ and we cover the rest of the grid with bricks.
A super-square is a square of bricks of side length $5\Delta^{2}R$ bricks. We need $150\Delta^{2}R$ to be slightly smaller than $T$. The might need a few extra rows and columns for divisibility reasons but for notational simplicity we ignore this difference and simply assume $T=150\Delta^{2}R$. Starting in the top left corner in each super-square we have $\Delta$ mini-squares along the diagonal. This leaves $\Delta^{2}R$ empty rows of bricks in the bottom of each super-square and $\Delta^{2}R$ empty columns at the right. By our placement of the designated survivor and the elimination of the top row and first column also the top left super-square looks essentially the same as other super-squares. Let us describe how to route horizontal paths.
[…]
For each $j$, we reserve $R$ bricks, all with even index, on the right perimeter of $s_{i}$ and the $k$th brick contains three paths that we route straight right to $c_{ij}^{k}$. Similarly we route paths straight left from the $k$th brick in the $i$th part of $s_{j}^{\prime}$ to the same column. This time using odd index bricks. The path is completed by using the suitable part of $c_{ij}^{k}$ to connect the two pieces.
Connecting mini-squares vertically is done completely analogously and we omit the description. Let us make a not very difficult observation.
Lemma 3.1
In each brick outside the mini-squares there is at most one vertical and one horizontal path.
Proof:  Let us look at the area to the right of the diagonal in one super-square and to the left of the diagonal in the super-square to its right. The only vertical parts of paths in this area are from the middle segments of the paths. By definition these are in disjoint columns.
The horizontal parts of paths in the this area are from either segment one or segment three. As the former are in even numbered rows of bricks and the latter in odd numbered rows these are also disjoint.
The argument for the area below a diagonal of one super-square and above the diagonal below, is completely analogous.
The remaining part is the area down to the right in each mini-square, i.e. in the intersection of columns with no mini-squares and rows with no mini-squares. In these bricks there are no paths.
We summarize the properties we need from this construction.
Lemma 3.2
The values of the path variables $y_{P}$ uniquely determines a matching on all paths and in any brick outside the mini-squares. In any mini-square except the designated survivor such that half of its adjacent $y_{P}$ variables are true we can find a unique matching of the remainder of this mini-square. In the designated survivor we can find a matching provided one more than half the adjacent variables is true.
3.2 Almost complete matchings and restrictions
We let an almost complete matching, usually denoted $\tau$, be an assignment to all variables $y_{P}$ such that exactly half the variables next to any mini-square are true. Note that many such $\tau$ do exist and in particular we can pick half the paths in any group of $2R$ paths to be of the each type. There are many other ways to pick $\tau$ but we do not need this explicitly.
Our restrictions give values to some variables $y_{P}$ and some remain unset. We rename the latter $z_{P}$ to distinguish new and old variables. While an individual $x_{e}$ might depend on many $y_{P}$ it only depends on at most three new variables.
Let us proceed to describe how to pick a random restriction from our space. We make uniformly random choices but if some choice is very unlucky we start all over again. Let $\pi_{1}$ be a fixed matching of all super-squares except that of the designated survivor in the top left corner. We pick a restriction $\sigma$ as follows.
  1. 1.
    Pick a uniformly random $\tau$ respecting the condition that half the variables next to any mini-square are true. If any group of $2R$ variables between two fixed mini-squares has fewer than $R/2$ variables of either value, restart. We denote this event as “lopsided group”.
  2. 2.
    Pick a random mini-square from each super-square, except the super-square of the designated survivor. Match these mini-squares according to $\pi_{1}$. Hopefully without causing confusion we keep the name $\pi_{1}$ for this matching of mini-squares. These nodes jointly with the designated survivor are the chosen mini-squares and we call this set $U$.
  3. 3.
    For each pair of mini-squares $(s_{1},s_{2})$ matched in $\pi_{1}$ pick one augmenting path of type 1 and convert it to type 0. These are called chosen paths. The choice is based on an advice string $B$ as defined below.
  4. 4.
    For each pair of mini-squares $(s_{1},s_{2})$ in $U$ in adjacent super-squares but not matched in $\pi_{1}$ pick one augmenting path of type 0 and make it a chosen path. The choice of the path is based on the advice string as defined below. This is also done with $s_{1}$ being the designated survivor.
Let us see that this defines a smaller instance of PHP in a set of new variables $z_{P}$ corresponding to the value of $y_{P}$ on all chosen paths.
3.3 The reduced instance
The nodes of the new instance are given by the elements in $U$. They naturally form a $(n/T)\times(n/T)$ grid. We have chosen paths that are of type 0 that connect any two adjacent elements in $U$ while for any other path $P$, the value of $y_{P}$ is now fixed.
By Lemma 3.2 if exactly one variable on a chosen path next to a mini-square in $U$ is true, then it is possible to find a matching of this mini-square. Thus the local conditions of a mini-square turn in to an axiom of the new instance of PHP. Let us see how to replace the old variables in a supposed proof with these new variables. First note that old variables not on edges in chosen mini-squares or in bricks with at least one chosen path are now fixed in a way respecting the corresponding axioms.
Consider a brick with at least one chosen path going through. There is only one chosen path between any two adjacent chosen mini-squares and it is not difficult to see that at most one chosen path goes through any brick. There are two possible matchings of this brick depending on the value of the corresponding variable $z_{P}$. If an edge $e$ is present in neither we replace $x_{e}$ by 0 and if it is present in both we replace it by 1. If $e$ is present in only one we replace it by $z_{P}$ or $\bar{z}_{P}$ in the natural way. It is easy to see that any axiom related to a node in the brick becomes true.
Similarly in a mini-square we have four (or fewer if it is on the perimeter) chosen paths next to it. These are controlled by four new variables that we here locally call $z_{i}$ for $1\leq i\leq 4$. The new local axiom is that exactly one of these four variables is true.
We have four different matchings of the mini-square depending on which $z_{i}$ is chosen to be true. Look at an edge, $e$ and suppose it appears in the matchings corresponding to $z_{2}$ and $z_{3}$ being true. In such a case we replace $x_{e}$ by $z_{2}\lor z_{3}$ and similarly in other cases. If $e$ is in none of the four matchings we replace $x_{e}$ by the constant 0 and if it is in all four we replace it by the constant 1.
It is easy to check that any original axiom inside the mini-square either reduces to true or that exactly one of the four $z_{i}$ is true. Indeed looking at the disjunctions replacing the four variables $x_{e}$ around any node, each $z_{i}$ appears in exactly one. We summarize the discussion of this section as follows.
Lemma 3.3
A restriction $\sigma$ reduces an axiom in the PHP in the grid of size $n\times n$ either to the constant true or an axiom in the PHP of the new variables $z_{P}$. Each variable $x_{e}$ is substituted by a conjunction of up to three literals.
Let us finally note that each variable $x_{e}$ that is not replaced by a constant or a single variable appears inside a chosen mini-square corresponding to a node, $v$, in the new instance. It is completely determined by the answer to the question “To which node is $v$ matched” and hence it is not very different from other variables. In particular since it is determined by one basic query there is no problem with a build-up of complicated expressions when we compose restrictions.
As in previous papers these full restrictions are not used in the main argument and we work with partial restrictions that we now turn to.
3.4 Partial restrictions
A full restriction is a very rigid object with exactly one live mini-square in each super-square and we proceed to make a more random looking object.
Let $k$ be a parameter equal to $C(n/T)^{2}\log n$ for a sufficiently large constant $C$. After we have completed the construction of $\sigma$ we add the following steps.
  1. 1.
    Pick, without replacement of mini-squares, $k$ uniformly random pairs of mini-squares in adjacent super-squares. These are picked one at the time and each is picked uniformly from the set of remaining possible pairs. This yields a matching $\pi_{2}$. If this pick is unbalanced as defined below redo this step.
  2. 2.
    Change the type of one augmenting path between any pairs of nodes in $\pi_{2}$ from 1 to 0. The choice of which of the at least $R/2$ paths of type 1 is based on the advice string and defined below.
The mini-squares picked during this process jointly with the chosen mini-square are considered “alive”. For any path, $P$ between two live mini-squares the corresponding value, $y_{P}$ is now considered undetermined while other $y_{P}$ variables are now fixed. It is not true that these undetermined variables can vary freely as in fact for each pair of live mini-square we may change the value of at most one of its adjacent paths from 0 to 1. No $y_{P}$ will be changed from 1 to 0. In spite of this we do not consider such a value to be known unless we have the full information of one of its end points.
As we pick $2k$ mini-squares we expect roughly [^3] The reason this is not exactly true is super-squares at the perimeter have only two or three neighboring super-squares. Such squares are less likely to have many live mini-squares. This results in a factor $(1+o(1))$ more mini-squares in other super-squares but this small factor does not matter and we ignore it. $2C\log n$ live mini-squares in any super-square. If this number is larger than $4C\log n$ for any super-square then we consider the pick unbalanced and we restart. For any two adjacent super-squares $S_{1}$ and $S_{2}$ we expect $C\log n/2$ pairs $(s_{1},s_{2})$ picked such that $s_{i}\in S_{i}$. If this number is smaller than $C\log n/4$ for any pair $(S_{1},S_{2})$ we consider this unbalanced and restart.
We call the resulting restriction $\rho$. It determines values for variables $y_{P}$ exactly as for $\sigma$. A variable $x_{e}$ is fixed to a constant unless it as influenced by at least one live mini-square. Either from being within such a mini-square or in a brick with at least one live path passing through.
3.5 Changing types of augmenting paths
In the above procedure, in two places we need to select an augmenting path and (possibly) change its type. This happens when changing a path from type 1 to type 0 because its end-points are matched in $\pi_{1}$ or $\pi_{2}$ and when opening up for changing the type from 0 to 1 by making it a chosen path.
We could accept to make this choice arbitrary by losing some factors of $R$ in our bounds, but as it is always nice to avoid unnecessary loss let us describe a more efficient choice.
The choices of the $2R$ variables in a group corresponds to a vector in $\{0,1\}^{2R}$ and we want to modify one coordinate in order to change the Hamming weight from $t$ to either $t+1$ or $t-1$ and let us suppose the latter. We want the choice to be limited and as invertible as possible. As the number of strings of weight $t$ and $t-1$ are different we cannot achieve perfect invertability. Suppose first that $t\leq R$.
Definition 3.4
Suppose $t\leq R$. A mapping $f$ mapping (${{2R}\choose t}$ to (${{2R}\choose{t-1}}$ and which maps each set to a subset, is a $k$-almost bijection if it is surjective and $|f^{-1}(x)|\leq k$ for any $x$.
The following below lemma is probably well known but as the proof is not difficult, we prove it. It is likely that 4 can be improved to 3 but this does not matter greatly for us, as this only affects unspecified constants.
Lemma 3.5
If $R/2\leq t\leq R$ then there is a 4-almost bijection.
Proof:  Consider a bipartite graph where the the left hand side elements are subsets of size $t-1$ and the right hand side elements are subsets of size $t$. Connect two sets iff one is a subset of the other. It is well known (see for instance Corollary 2.4 in [4]) that this graph has a matching, $M_{1}$, of size ($2R\choose t-1$.
Modify the construction by making three copies of each left hand side node. Each copy is again connected to any set that contains it. It follows by the LYM inequality (stated as Theorem 3.3 in [4]) that this graph has a matching $M_{2}$ of size ($2R\choose t$.
Now define $f(x)$ as follows. If $x$ is matched in $M_{1}$ let it be its partner in this matching. If $x$ is not matched in $M_{1}$ define $f(x)$ to be the partner under $M_{2}$.
Due to the first condition $f$ is onto. The property that $|f^{-1}(y)|\leq 4$ for follows as a preimage of $y$ is either its partner under $M_{1}$ or a partner of one of its three copies under $M_{2}$.
Taking the complement of both input and output we define a $4$-almost bijection, $g$ mapping ($2R\choose t$ to ($2R\choose{t+1}$ for $R\leq t\leq 3R/4$. We use $f$ and $g$ to guide our choices and in addition we have two bits of advice for each group.
Definition 3.6
For each group of $R$ paths we have two bits in the advice.
When we want to convert a path from type 1 to type 0 between $s_{i}$ and $s_{j}^{\prime}$ we look the types of all paths between the two mini-squares. This is a vector, $v$, in $\{0,1\}^{2R}$ which by the non-lopsidedness has Hamming weight $t$ which is in the interval $[R/2,3R/2]$. If $t>R$ we look at $g^{-1}(v)$ and consider the two bits of advice, $b_{1}$ and $b_{2}$. All we need to do is to ensure that each choice in $g^{-1}(v)$ is possible but to be explicit we can proceed as follows.
  • If $g^{-1}(v)$ is of size one we pick the unique element.
  • If $g^{-1}(v)$ is of size two we use $b_{1}$ to make the choice.
  • If $g^{-1}(v)$ is of size three then $b_{1}=b_{2}$ we pick the lexicographically first path and otherwise we use $b_{1}$ to choose between the other two paths.
  • If $g^{-1}(v)$ is of size four then use the pair $(b_{1},b_{2})$ to make the choice.
The reason for the advice string is to get a pure counting argument when we later analyze probabilities. It would have worked to pick a random element from $g^{-1}(v)$.
To make the situation uniform we have two advice bits for any pair of mini-squares in adjacent super-squares. Thus most of these bits are never used. We let $B$ denote the values of all these bits.
If $t\leq R$ we instead consider $f(v)$ and change the type of the corresponding path. Finally if we want to allow to change the weight from $t$ to $t+1$ we reverse the two cases.
3.6 Analyzing the probability of a restart
We make a restart either because of a lopsided group or an unbalanced pick of $\pi_{2}$ and we analyze these separately. We start with lopsided groups.
Lemma 3.7
The probability that uniformly random $\tau$ has lopsided group is $O(n^{2}2^{-cR})$ for a positive constant $c$.
Based on this lemma we fix $R$ to be $C\log n$ for a sufficiently large constant $C$ such that the probability of having a lopsided group is $o(1)$. Let us prove Lemma 3.7.
Proof:  The almost complete matching $\tau$ is defined by the variables $y_{P}$ which we in this section choose to take values 1 and $-1$. For a uniformly random assignment to all variables, let $Z$ be the vector of all mini-square sums. Let us denote the number of mini-squares by $d$ making $Z$ an integer vector of length $d$. When constructing $\tau$ we are conditioning on the event $Z=0^{d}$.
Fix any two mini-squares $s_{1}$ and $s_{2}$ and let $g$ be the group of $2R$ variables associated with paths between $s_{1}$ and $s_{2}$. Let $X_{g}$ be sum of the variables in this group. We want to estimate the probability that $X_{g}=x$ where $x$ is either at least $R$ or at most $-R$. Let $Z^{\prime}_{g}$ be the set of mini-square sums when the paths between $s_{1}$ and $s_{2}$ are removed. As these two mini-squares also have other adjacent augmenting paths this is still a vector length $d$. Let $v_{x}$ be the vector of length $d$ that has $x$ at positions $s_{1}$ and $s_{2}$ and is otherwise 0. We want to estimate
$Pr[X_{g}=x\ |\ Z=0^{d}]=Pr[X_{g}=x\land Z=0^{d}]/Pr[Z=0^{d}]$
which equals
$Pr[X_{g}=x\land Z_{g}^{\prime}=-v_{x}]/Pr[Z=0^{d}]$
and as the two events are independent this equals
$Pr[X_{g}=x]Pr[Z_{g}^{\prime}=-v_{x}]/Pr[Z=0^{d}].$
We have the following lemma of which we postpone the proof.
Lemma 3.8
For any outcome $v\in\mathbb{Z}^{d}$ we have $Pr[Z_{g}^{\prime}=v]\leq Pr[Z_{g}^{\prime}=0^{d}]$. The similar statement is true for any multi-graph where each edge appear an even number of times.
In view of the lemma we get the upper bound
$Pr[X_{g}=x]Pr[Z_{g}^{\prime}=0^{d}]/Pr[Z=0^{d}]$
for the probability we want to estimate. Clearly $Pr[Z=0^{d}]\geq Pr[X_{g}=0]Pr[Z_{g}^{\prime}=0^{d}]$ and substituting this into the equation we get the upper bound $Pr[X_{g}=x]/Pr[X_{g}=0]$. When $|x|\geq R$ this probability is $2^{-cR}$ for some explicit $c$ and since there are at most $n^{2}$ pairs of mini-squares the lemma follows.
Let us prove Lemma 3.8.
Proof:  Let $f(v)$ be the probability that $Z^{\prime}_{g}=v$. As $v$ is the vector sum of contributions of single $y_{P}$, $f(v)$ is a giant convolution. A typical term, $f_{0}$, is a probability distribution that give weight $\frac{1}{2}$ to the vector $v_{1}$ and weight $\frac{1}{2}$ to the vector $-v_{1}$ where $v_{1}$ is similar to the vector $v_{x}$ in the above proof. Consider the corresponding Fourier expansion
$\hat{f}_{0}(x)=\sum_{v}f_{0}(v)e^{2\pi i(v,x)}$
where $(v,x)$ is the inner product of $v$ and $x$ and $x$ belongs to the $d$-dimensional torus. As $f_{0}$ is symmetric under negation, $\hat{f}_{0}$ is real-valued. Summing all variables in a given group gives a distribution function which is a $2R$-fold convolution of $f_{0}$. Its Fourier expansion represents $\hat{f}_{0}^{2R}$ which is a positive function. Summing the contribution from all edges corresponds, on the Fourier side, of taking the product. We conclude that
$\hat{f}(x)=\sum_{v}f(v)e^{2\pi i(v,x)}$
is a real-valued positive function. The largest Fourier coefficients of such a function is the constant coefficient and this is exactly what we wanted to prove.
Let us next discuss the balance condition when picking $\pi_{2}$.
Lemma 3.9
The probability that $\pi_{2}$ is unbalanced is $O(n^{-2})$ provided $C>C_{0}$ for some fixed constant $C_{0}$.
Proof:  As this is a very standard argument let us only sketch it. If we picked the pairs of mini-squares with replacement the lemma would be completely standard. Let us analyze the dynamic process. To see that we do not pick more than $4C\log n$ mini-squares in any super-square with high probability we note two facts.
[…]
  • In view of this the probability that any single mini-square is picked is only a $(1+o(1))$ factor larger compared to the procedure with replacement.
That we are unlikely to pick many mini-squares in a single super-square now follows from the corresponding result for the process of picking with replacement.
We turn to the condition that we have at least $C\log n/4$ pairs in any two adjacent super-squares. Also this analysis is completely standard if edges are picked with replacement. If we condition on not picking more than $4C\log n$ mini-squares in any super-square the probability that a picked edge is between two given super-squares does not decrease by more than a factor $1-o(1)$. Hence the probability of picking very few edges between two given super-squares in the process without replacement is not so different compared to the probability of the same event in the process with replacement. We leave it to the reader to fill in the details.
4 The switching lemma
In this section we establish the following basic switching lemma.
Lemma 4.1
There is a constant $A$ such that the following holds. Suppose there is a $t$-evaluation that includes $F_{i},1\leq i\leq m$ in its range and let $F=\lor_{i=1}^{m}F_{i}$. Let $\sigma$ be a random restriction from the space of restrictions defined in Section 3. Then the probability that $F\lceil_{\sigma}$ cannot be represented by a decision tree of depth at most $2s$ is at most
$\Delta(A(\log n)^{3}t\Delta^{-1})^{s}.$
Proof:  We are interested in a $\sigma$ that gives a long path in the decision tree. As in previous papers [10], [11] we explore what has been called “the canonical decision tree” under the partial restriction $\rho$ which we from now on call simply “a restriction” dropping the word “partial”.
For any variable $x_{e}$ we define its influential mini-square(s). This is either a single mini-squares or two mini-squares. We want the property that if we know the values of all $y_{P}$ around the influential mini-square(s) then this uniquely determines the value of $x_{e}$. If $e$ is within a mini-square then this mini-square is its influential mini-square(s). If $e$ is in a brick outside the mini-squares then the influential mini-squares are the closest end point(s) of the live path(s) in this brick. By “closest” we here mean a live mini-square such that if we know the value of all adjacent variables $y_{P}$ then we know the value of the $x_{e}$. We do not consider the value of $x_{e}$ known unless we completely know the situation at its influential mini-square(s). If such a mini-square is alive the extra information needed is provided by something we call a matched pair.
Definition 4.2
A matched pair $(s_{1},s_{2})$ is the information that $s_{1}$ and $s_{2}$ should be matched by changing the type of an augmenting path from 0 to 1 between the two mini-squares.
We use this concept for pairs of neighboring chosen mini-squares and for pairs in the matching $\pi_{2}$. In both cases the identity of the path to change follows from the context. For pairs of chosen mini-squares it is the chosen path and for pairs in $\pi_{2}$ it is the path the was modified when it was first picked.
A matched pair $(s_{1},s_{2})$ fixes, once and for all, the values of all variables $y_{P}$ where $P$ is a path with an end-point $s_{1}$ or $s_{2}$
We now proceed to define the canonical decision tree. The process is guided by $\rho$ and a set $I$ of matched pairs. The support of the set $I$ is the mini-squares appearing in any matched pair and this is initially empty.
Letting $T_{i}$ denote $\varphi(F_{i})$, we go over the branches of $T_{i}$ for increasing values of $i$ and in a fixed order for each tree. The forceable branch is the first branch leading to a one that can be followed given the current $I$ and $\rho$. Let us be formal.
Before stage $j$ we have information set $I^{j}$ in the form of some matched pairs. It contains some pairs from $\pi_{2}$ and some pairs based on answers in the decision tree. In stage $j$ we have forcing information $J_{j}$ that makes us follow the forceable branch and it contains.
  1. 1.
    A set of matched pairs from $\pi_{2}$.
  2. 2.
    A set of matched pairs of chosen mini-squares, consistent with the information set $I^{j}$.
By “consistent” we mean that the resulting partial matching on chosen mini-squares is locally consistent in the sense of Definition 2.1.
For any chosen mini-square mentioned in $J_{j}$ we ask for its partner in the decision tree. This information, jointly with the matched pairs from $\pi_{2}$ in $J_{j}$, forms the $j$th information set, $I_{j}$, and we set $I^{j+1}=I^{j}\cup I_{j}$.
Given $I^{j+1}$ we can determine whether the forceable branch is followed. If it is, we answer 1 in the canonical decision tree and halt the process. Otherwise we go to the next stage and look for the next forceable branch. If there is no more forceable branch we halt with answer 0. Let $T$ be the resulting decision tree. To see that this is a an acceptable choice for $\varphi(F)$, we have a pair of lemmas.
Lemma 4.3
Let $\gamma$ be a set of answers in the decision tree. If there is no forceable branch given this information, then, for each $i$, $T_{i}\lceil_{\sigma\gamma}$ is a 0-tree.
Proof:  Suppose there is a locally consistent branch in $T_{i}\lceil_{\sigma\gamma}$ that leads to a leaf labeled one. The information used to follow this branch can be used as forcing information.
Lemma 4.4
Let $\gamma$ be a set of answers in the decision tree. If we answer 1 then there is an $i$ such that $T_{i}\lceil_{\sigma\gamma}$ is a 1-tree.
Proof:  In fact for $T_{i}$ used for the construction of the forceable branch we have now reached a leaf that is labeled 1.
We use Razborov’s labeling argument [18] to analyze the probability that we make $2s$ queries in the canonical decision tree. Take any branch of this length and suppose it was constructed during $g$ stages using information sets $J_{j}$. Let $J^{*}=\cup_{j=1}^{g}J_{j}$ and as any query is a result of including an element in $J^{*}$ we know that it contains at least $s$ pairs and for notational convenience we assume that this number is exactly $s$. Note that $J^{*}$ may not be locally consistent when seen as a partial matching on the chosen nodes, but this is not required. We proceed to analyze the probability that the process results in a $J^{*}$ of this size. Let us start with an easy observation.
Lemma 4.5
The support sets of $J_{j}$ are disjoint. The support of $J_{j}$ is also disjoint with the support of $I_{i}$ as long as $i\not=j$.
Proof:  The parts coming from $\pi_{2}$ are clearly disjoint as $\pi_{2}$ is a matching. The pairs of chosen nodes are also disjoint as any mini-square included in $J_{j}$ is included in $I_{j}$ and later $J_{j^{\prime}}$ are disjoint from $I_{j}$.
In the spirit of [10] and [11] we want to find a restriction $\rho^{*}$ with fewer live mini-squares and then show how to reconstruct $\rho$ from $\rho^{*}$ and some external information. First note that it is possible for several tuples $(\tau,U,\pi_{2},B)$ to produce the same $\rho$. For this reason we construct a quadruple $\tau^{*},U^{*},\pi_{2}^{*},B^{*}$ and use this together with external information to reconstruct $\tau,U,\pi_{2},B$. The quadruple $\tau^{*},U^{*},\pi_{2}^{*},B^{*}$ corresponds to a restriction $\rho^{*}$.
Informally we want $\rho^{*}$ to be $\rho$ modified by the information set $J^{*}$. In other words for a matched pair in $J^{*}$ we change the type of one augmenting path between the two mini-squares and now consider the two endpoints to be dead. We proceed to find a quadruple that gives this restriction.
As a first approximation we let $\pi_{2}^{*}$ be $\pi_{2}$ with all matched pairs in $J^{*}$ removed and $U^{*}=U$. We need to process the matching pairs, $(s_{1},s_{2})$ of chosen nodes in $J^{*}$. We first remove these two mini-squares from $U^{*}$. If there is a pair $(s_{1}^{\prime},s_{2}^{\prime})$ in $\pi_{2}^{*}$ where $s_{i}^{\prime}$ is in the same super-square as $s_{i}$ we remove this pair from $\pi_{2}^{*}$ and put $s_{1}^{\prime}$ and $s_{2}^{\prime}$ in to $U^{*}$. If there is no such edge we allow for two “holes” of two empty super-squares in $U^{*}$.
After we have constructed $U^{*}$ and $\pi^{*}_{2}$ we now construct $\tau^{*}$ and $B^{*}$ to complete the quadruple. Let us start by setting $\tau^{*}$ to take the values given by $\rho$.
Let us, again in ±$\pm 1$ sum up the number the neighboring variables $y_{P}$ around any mini-square. Due to the changing of the variables corresponding to $\pi_{1}$ and $\pi_{2}$ the sum for any live mini-square, except the designated survivor is $-2$ while the sum around the designated survivor is 0. The sum around the dead mini-squares is 0.
For pairs from $\pi_{1}$ or $\pi_{2}$ that remains matched either in $\pi_{2}^{*}$ or $\pi_{1}^{*}$ we restore the value of $\tau^{*}$ to take the value of $\tau$ and we let the corresponding bits of $B^{*}$ equal to those of $B$.
For pairs in $J^{*}\cap\pi_{2}$ we change the value of the paths used in its construction from 0 to 1 in $\tau^{*}$. We set the bits $B^{*}$ to the value in $B$.
For the pairs of chosen mini-squares in $J^{*}$ we change the value of the $y_{P}$ corresponding to the chosen path from 00 to $1$, we set the values of $B^{*}$ equal to the bits of $B$.
For mini-squares in $U^{*}$ that either are new (because they came from $\pi_{2}$) or has lost their partner (due to the partner belonging to a pair in $J^{*}$) we proceed as follows. If it has a partner in $\pi^{*}$ pick one possible path $P$ between it and its partner and change the value of $y_{P}$ from 0 to 1 in $\tau^{*}$. Set the bits of $B^{*}$ so that this would have been the path picked by $\pi_{1}$.
We have created an almost complete matching with some exception. If the designated survivor belonged to a pair of $J^{*}$ its sum is now +2.
Any mini-square in $U^{*}$ without a partner in $\pi^{*}_{1}$ has sum -2. This can either be the replacement of the designated survivor in its square or due to a hole in $U^{*}$.
We continue to define a process that, using external information, reconstructs $\tau$, $U$, $\pi_{2}$ and $B$ from $\tau^{*}$, $U^{*}$, $\pi_{2}^{*}$ and $B^{*}$ and the trees $T_{i}$. The “star quadruple” determines the values of of $y_{P}$ in the same way as the non-stared quadruple determined values for $y_{P}$. To be more precise the process is the following.
  • Start with values given by $\tau^{*}$.
  • Define chosen mini-squares by $U^{*}$.
  • For any pair of nodes in $\pi_{1}$ (as a matching of super-squares) find path connecting this pair as indicated by $B^{*}$. Change the corresponding value from $1$ to 00.
  • For any pair in $\pi_{2}^{*}$ find one path using $B^{*}$ and change the value from 1 to 0.
  • Let the live mini-squares be the elements of $U^{*}$ and the mini-squares of $\pi_{2}^{*}$.
  • Consider a value $y_{P}$ to be fixed unless if it is between two live mini-squares.
Call this restriction $\rho^{*}$. Let us first the key fact.
Lemma 4.6
The restriction $\rho^{*}$ forces the first forceable branch to be followed.
Proof:  We claim that any variable $y_{p}$ forced by $\rho$ is also forced by $\rho^{*}$ and to the same value. This follows as any such variable requires all its influential mini-squares to be dead. Any mini-square that this is dead in $\rho$ is also dead in $\rho^{*}$. As the situation around it has not changed it is forced to the same value.
The additional information to follow the first forceable branch is given by $J_{1}$. As this information is included in $\rho^{*}$ the first forceable branch is followed.
To correctly identify the first forceable branch is an important step in reconstruction. Unfortunately there might be earlier paths forced to one by $\rho^{*}$. The reason is that the information that makes us follow such a path might come from several different $J_{i}$ and there is no guarantee that these matchings are locally consistent and hence it might not constitute an acceptable set $J_{1}$.
The optimistic view is that any such path still would help us as it must point to some live mini-squares. We are, however, not able to use this and we need to make sure that the reconstruction is not confused. We introduce the concept of “signature” to enable us to identify the first forceable branch.
Definition 4.7
The signature of a live mini-square determines whether it is chosen and in such a case which direction it is matched in its forceable branch.
The reconstruction is now as follows. We start with $\rho^{*}$ and we reconstruct $I_{i}$ and $J_{i}$ in order. We let $\rho_{j}^{*}$ be the restriction obtained from $\rho$ jointly with $I_{i}$ for $i<j$ and $J_{i}$ for $i\geq j$. Thus $\rho^{*}_{1}$ is simply $\rho^{*}$ which is the starting point. We have a set $E$ of prematurely found chosen mini-squares jointly with their signatures. This is initially the empty set. We proceed as follows.
  1. 1.
    Find the next branch forced to one in any of the $T_{i}$ by $\rho_{j}^{*}$.
  2. 2.
    Find, if any, mini-square of $E$ whose information is used to follow this path. If this information is not locally consistent with $I^{j-1}$, go to the next branch.
  3. 3.
    Read a bit to determine whether there are more live variables to be found on this branch. In such a case, read a number in $[t]$ to determine its position. If this variable has two influential mini-squares, read another two bits to determine which of these two mini-squares are alive. For any alive influential mini-square, retrieve the corresponding signature(s) from external information and, if chosen, include the mini-squares in $E$. Go to step 2.
  4. 4.
    Reconstruct $I_{j}$ and $J_{j}$. Details below.
  5. 5.
    Remove any chosen mini-square included in $J_{j}$ from $E$.
We need a lemma.
Lemma 4.8
If the information from $E$ used on the branch leading to a one is consistent with the information in $I^{j-1}$ then it is the forceable branch.
Proof:  Suppose $v\in E$, which by definition, is a chosen mini-square, and was included in $J_{j^{\prime}}$ for some $j^{\prime}\geq j$. As the current path is forced to one it must have been a potential forceable branch. If it was not the actual forceable branch it must be that at least one of the matched pairs needed to follow this branch was not allowed. Forcing information from $\pi_{2}$ is always allowed and thus the only problem could have been consistency on the chosen mini-squares. If none of the elements used on the current path gives any conflict with $I^{j-1}$, it was allowed and hence it must be the forceable branch.
We need to discuss how to reconstruct $I_{j}$ and $J_{j}$ and start with the latter. For each matched pair in $J_{j}$ we have recovered at least one end-point from the forceable branch. If needed we use external information to recover the other end-point. This costs at most $\Delta$. Whenever we recover a pair of adjacent mini-squares we use external information to recover the advice bits. This only costs 2 bits and thus we below focus on identifying mini-squares and do not mention the reconstruction of advice bits.
For each element in $J_{j}$ we need to discover whether it is chosen (this is one bit of external information) and to which node is is matched in $I_{j}$. If it is not chosen then the information in $I_{j}$ is the same as the one in $J_{j}$. If it is chosen then the partner is either in $J_{j}$ and easy to find or it is live in $\rho^{*}_{j}$. This follows as if it is chosen and belongs to $I_{j}$ but not $J_{j}$ then, by Lemma 4.5 it cannot belong to $J^{*}$. In the case when it is alive we can, at cost $2C\log n$, find its identity. Once we have identified the two partners in a matching, at cost $O(1)$ we can reconstruct which augmenting path was used.
We reconstruct $I_{j}$ and $J_{j}$ and compute $\rho_{j+1}^{*}$ and proceed to the next stage. At the end of this process we have reconstructed $\rho$ and in the process we have also identified all the sets $I_{j}$ and $J_{j}$ and we know which pairs in $J_{j}$ are chosen and which belong to $\pi_{2}$. For the pairs of chosen mini-squares we put them back in to $U$ and any replacements are moved back to $\pi_{2}$. At cost $O(1)$ we can identify the advice bits of any single mini-square involved in a move. This way we recover $U$, the full $\pi_{2}$, and $B$. Once we have these we can read off the original $\tau$.
Let us calculate how much information was used. For each edge in $J^{*}$ at least one end-point is discovered at cost at most $t$ and the other at cost at most $\Delta$. On top of this we have information such as the whether it is chosen and the signature at cost $O(1)$. This gives a total cost of at most $2^{O(s)}(t\Delta)^{s}$. In reconstructing $I_{i}$ we might incur an additional cost of $O(\log n)$ per mini-square for an additional factor of $2^{O(s)}(\log n)^{2s}$ and finally a factor $2^{O(s)}$ to get the advice bits for nodes that move out and into $U$. Thus the total cost for the external information is bounded by $2^{O(s)}(\log n)^{2s}(t\Delta)^{s}$.
We proceed to compare the number of quadruples $\tau,U,\pi_{2},B$ to the number of quadruples $\tau^{*},U^{*},\pi_{2}^{*},B^{*}$, and let us first assume that there are no holes. Both $U$ and $U^{*}$ contain one mini-square from each super-square but there is a difference that $U$ contains the designated survivor while $U^{*}$ might contain a different mini-square from this super-square. Thus the number of possibilities for $U^{*}$ is a factor $\Delta$ larger than the number of possibilities for $U$.
Both $\tau$ and $\tau^{*}$ are essentially almost complete matchings. We do not allow $\tau$ to be lopsided but this only reduces the number of possibilities, by Lemma 3.7, by a factor $1+o(1)$. It is the case that $\tau^{*}$ is only slightly more general but we only need an upper bound on the number of possibilities. On the other hand if $U^{*}$ does not contain the designated survivor there are two mini-squares in $\tau^{*}$ where the sum is not zero. This is the designated survivor and its replacement in $U^{*}$ in the same super-square. The number of $\tau^{*}$ with these fixed sums is, however, smaller by Lemma 3.8. There is no difference between $B$ and $B^{*}$ and thus the number of alternatives is the same. We conclude that if we denote the number of different $\tau,U$ and $B$ by $N$, then the number of triples $\tau^{*},U^{*}$, and $B^{*}$ is bounded by $\Delta N(1+o(1))$.
The big difference is in the probability of picking the restriction $\pi_{2}$ and $\pi_{2}^{*}$. Let $D$ denote the total number of possible pairs of mini-squares. We have that $D=\Delta^{2}(n/T)^{2}(2+o(1))$. This follow as there are $\Delta(n/T)^{2}$ mini-squares. Most of them (except at the perimeter) have $4\Delta$ possible partners and hence the number of pairs is $(1+o(1))\Delta(n/T)^{2}4\Delta/2$. We have the following lemma.
Lemma 4.9
Assuming $\Delta=\omega(\log n)$ then $Pr[\pi_{2}]\leq 2^{o(s)}(k/D)^{s}Pr[\pi_{2}^{*}]$ which in its turn is bounded by $2^{O(s)}(\log n/\Delta^{2})^{s}Pr[\pi_{2}^{*}]$.
Proof:  The big difference between the two is that $\pi_{2}$ is of size $k$ while $\pi_{2}^{*}$ is of size only $k-s$. The former is also not unbalanced, but as this, by Lemma 3.9, only changes the probability by a factor $(1+o(1))$ and we can let this factor be part of the error term. The way we define the probability space of partial matchings is slightly unusual and we need to be careful. There are $k!$ different orders to pick $\pi_{2}$. The probability of picking $\pi_{2}$ in a particular order $\gamma$ is $\prod_{i=1}^{k}t_{i,\gamma}^{-1}$ where $t_{i,\gamma}$ is the number of pairs of mini-squares available at stage $i$ assuming the order $\gamma$.
Similarly the probability of $\pi_{2}^{*}$ can be written as
$\sum_{\gamma^{*}}\prod_{i^{*}=1}^{k-s}t_{i^{*},\gamma^{*}}^{-1},$
where $\gamma^{*}$ determines the order in which its $k-s$ pairs are added. Each picked pair of mini-squares eliminates at most $8\Delta$ possible other pairs and hence each $t_{i,\gamma}$ and $t_{i^{*},\gamma^{*}}$ is at least $D-8k\Delta$. The product giving the probabilities do have many factors so we need more precise information.
Say that $\gamma$ and $\gamma^{*}$ are associated (written as $\gamma\sim\gamma^{*}$) if the $k-s$ common elements are in the same order. In such a situation for each $i^{*}$ there is an $i$ corresponding to the same element such that $t_{i^{*},\gamma^{*}}\geq t_{i,\gamma}\geq t_{i^{*},\gamma^{*}}-8s\Delta$. Restricting the product to these paired indices and using $t_{i,\gamma}\geq D/2$ we have
$\prod_{i}t_{i,\gamma}\geq(1-16s\Delta/D)^{k}\prod_{i^{*}}t_{i^{*},\gamma^{*}}.$
As $k=\Theta(\log nD/\Delta^{2})$ we have that
$(1-16s\Delta/D)^{k}=2^{-\Theta((s\log n)/\Delta)}=2^{o(s)},$
given the assumption on $\Delta$. As each factor omitted is at least $D-8k\Delta$ we can conclude that
$\prod_{i}t_{i,\gamma}\geq D^{s}2^{-o(s)}\prod_{i^{*}}t_{i^{*},\gamma^{*}}.$
There are $k!/(k-s)!\leq k^{s}$ different $\gamma$ associated with each $\gamma^{*}$. We conclude that
$Pr[\pi_{2}]$ $=$ $\sum_{\gamma}\prod_{i=1}^{k}t_{i,\gamma}^{-1}=\sum_{\gamma^{*}}\sum_{\gamma\sim\gamma^{*}}\prod_{i=1}^{k}t_{i,\gamma}^{-1}\leq\sum_{\gamma^{*}}\sum_{\gamma\sim\gamma^{*}}D^{-s}2^{o(s)}\prod_{i^{*}=1}^{k-s}(t^{*})_{i^{*},\gamma^{*}}^{-1}\leq$
$\leq$ $\sum_{\gamma^{*}}(k/D)^{s}2^{o(s)}\prod_{i=1}^{k-s}(t)_{i^{*},\gamma^{*}}^{-1}\leq(k/D)^{s}2^{o(s)}Pr[\pi_{2}^{*}],$
giving the bound of the lemma. The second bound is obtained by using the definitions of $k$ and $D$.
Let us finish the proof of Lemma 4.1. We can produce a random tuple $(U,\tau,B,\pi_{2})$ that forces a branch of length at least $s$ in the following way.
  • Pick a random $(U^{*},\tau^{*},B^{*},\pi^{*}_{2})$.
  • Pick external information and use the above inversion process to obtain $(U,\tau,B,\pi_{2})$.
Up to a factor $(1+o(1)$ we have $\Delta N$ possibilities for $U^{*},\tau^{*},B$ and $N$ possibilities for $U,\tau,B$ and the uniform distribution on both spaces. If follows from Lemma 4.9 that the probability of the produced four-tuple is at most $\Delta 2^{O(s)}(\log n/\Delta^{2})^{s}$ that of initial four-tuple.
Recalling that we have $2^{O(s)}(\Delta t)^{s}(\log n)^{2s}$ possibilities for the external information, we can conclude that if we sum $\Pr[(U,\tau,B,\pi_{2})]$ over all tuples producing a canonical decision tree of depth at least $s$, then this is bounded by
$2^{O(s)}(\Delta t)^{s}(\log n)^{2s}\Delta(\log n/\Delta^{2})^{s}\leq 2^{O(s)}((\log n)^{2}t/\Delta)^{s}.$
This proves Lemma 4.1 in the case when there are no holes.
If we allow $r$ holes then the number of possibilities of $U^{*}$ increases by a factor at most $n^{2r}$. The number of $\tau^{*}$ is different as some sums are no longer 0, but the sum at each mini-square is uniquely determined by $U^{*}$. Once this is fixed the number of $\tau^{*}$ is at most the number of almost complete matchings by Lemma 3.8.
For there to be a hole, by the balance condition, all the at least $C\log n/8$ edges between two super-squares must be present in $J^{*}$. This implies that $r=O(s/\log n)$ and the factor $n^{O(r)}$ can be absorbed in the factor $2^{O(s)}$. This completes the proof also in this case.
5 The lower bound for proof size
In this section we establish one of the two main theorems of the paper.
Theorem 5.1
Assume that $d\leq O(\frac{\log n}{\log{\log n}})$ and let $n$ be an odd integer. Then any depth-$d$ Frege proof of the functional onto PHP on the $n\times n$ grid requires total size
$\mbox{exp}(\Omega(n^{1/(2d-1)}(\log n)^{O(1)})).$
Proof:  The proof follows the standard path. We use a sequence of restrictions and after the $i$th restriction any sub-formula of the proof of original depth at most $i$ is in the range of the $t$-evaluation.
Assuming this, after $d$ restrictions, any sub-formula of the proof is in the range of the $t$-evaluation. By Lemma 3.3 what remains is a smaller functional onto PHP instance and by Lemma 2.14, provided the size of the remaining grid is significantly larger than $t$, the proof cannot derive contradiction. Let us give some details.
In the base case of depth 0, each literal naturally gives a 1-evaluation. Now suppose that size (and hence the number of sub-formulas) of the proof is $2^{S}$.
We apply a sequence of restrictions, the first one with $\Delta=\Omega((\log n)^{3})$ and later with $\Delta=\Omega(S(\log n)^{3})$ where we assume that the implied constant is large enough. We claim that, with high probability, after step $i$ each sub-formula that was originally of depth at most $i$ is now in the range of a $2S$-evaluation.
Indeed, consider Lemma 4.1 and take constants large enough so that the failure probability for an individual formula is bounded by $2^{-2S}$. By the union bound, except with probability $2^{-S}$, at step $i$ all formulas originally of depth $i$ are in the range of of a $2S$ evaluation.
The parameter $n$ turns into $\Theta(n/R\Delta^{2})$ by one application of a single restriction and hence after the $d$ restrictions, and using that $R=\Theta(\log n)$, we are left with a grid of size
$n2^{-cd}S^{2-2d}(\log n)^{-7d}$
for some constant $c$. There is a $2S$-evaluation which have every sub-formula of the proof in its range. By Lemma 2.14 provided that
$n\geq 300\cdot 2^{cd}(\log n)^{7d}S^{2d-1}.$
such a proof cannot refute the PHP. This concludes the proof of the theorem.
Next we turn to the multi-switching lemma.
6 Multi-switching
The extension to multi-switching only requires minor modifications and we start by stating the lemma.
Lemma 6.1
There is a constants $A$ such that the following holds. Consider formulas $F_{i}^{m}$, for $m\in[M]$ and $i\in[n_{m}]$, each associated with a decision tree of depth at most $t$ and let $F^{m}=\lor_{i=1}^{n_{m}}F_{i}^{m}$. Let $\sigma$ be a random full restriction from the space of restrictions defined in Section 3. Then the probability that $(F^{m})_{m=1}^{M}$ cannot be represented by an $\ell$ common partial decision tree of depth at most $4s$ is at most
$\Delta M^{4s/\ell}\big{(}A(\log n)^{7}t\Delta^{-1}\big{)}^{s}.$
Proof:  As in the proof of the single switching lemma we outline a process to construct an $\ell$ common decision tree. We prove that this process is unlikely to ask more than $4s$ variables in a similar way to the single switching lemma.
Let us define the process of constructing a $\ell$ common decision tree. We treat the formulas for increasing values of $m$. We have counter $j$, starting at 0, indicating the number of times we have found a long branch in a decision tree. We have an information set $I^{+}$ which initially is empty.
If a formula $F^{m}$ admits a decision tree of depth at most $\ell$ under $\rho$ and the current $I^{+}$ then we proceed to the next formula. If not, we set $m_{j}=m$ and execute the process of forming a canonical decision tree. It discovers a branch of length at least $\ell$ in this tree and constructs the corresponding sets $J_{i}^{j}$ for $i=1,2\ldots g_{j}$. We also have the corresponding information sets $I_{i}^{j}$.
After this branch has been discovered we ask, now in the $\ell$ common decision tree, for the partner of any chosen element in $I^{j*}=\cup_{i=1}^{g_{j}}I_{i}^{j}$. The answers jointly with the matching pairs from $\pi_{2}$ in $I^{j*}$ form a new information set $I^{j*+}$. This information set is added to $I^{+}$ and we look for the next formula that needs a decision tree of depth at least $\ell$. It might be the same formula $F^{m_{j}}$, but the situation is new as we an updated set $I^{+}$ and we know that the previous forceable branch is now not followed.
Properties follow along the same lines as in the single switching lemma and we start with the lemma corresponding to Lemma 4.5.
Lemma 6.2
The support sets of $J_{i}^{j}$ are disjoint. The support of $J_{i}^{j}$ is also disjoint with the support of $I_{i^{\prime}}^{j^{\prime}}$ as long as $i^{\prime}\not=i$ or $j^{\prime}\not=j$.
Proof:  The parts from $\pi_{2}$ are obviously disjoint and we need only consider the chosen mini-squares. For the same value of $j$, this lemma follows by the same argument as Lemma 4.5. As $I^{j*+}$ contains each $J^{j}_{i}$ and sets $J_{i}^{j^{\prime}}$ for $j^{\prime}>j$ are disjoint from $I^{j*+}$ the general case is also true.
Let $J^{j*}=\cup_{i}J_{i}^{j}$ be the information set obtained during the $j$th stage. It contains at least $\ell$ matched pairs and let us assume that the true number is $s_{j}$. Let $J^{*}=\cup_{j}J^{j*}$ and define $\rho^{*}$ as $\rho$ joint with information set $J^{*}$. We need to describe a process to find $\rho$ from $\rho^{*}$ and the set of formulas.
Let us immediately point out that the information which $F^{m_{j}}$ are processed is contained in the external information. As at most $4s/\ell$ formulas are processed this gives an extra factor $M^{4s/\ell}$ in the size of the external information. This factor appears in the bounds of the lemma and hence we proceed to analyze the rest of the process.
Once we know the value of $m_{j}$ we can run the reconstruction process based on $F^{m_{j}}$ finding the sets $I_{i}^{j}$ and $J_{i}^{j}$ exactly as in the single switching lemma. There is no need to change this procedure.
A difference compared to the single switching lemma is the sets $I^{*j+}$. Again the information pieces from $\pi_{2}$ do not give a problem as they never change. For each chosen mini-square in $I_{i}^{j}$ we need to specify a partner inside $I^{*j+}$. This may or may not belong to some $I_{i^{\prime}}^{j}$. If it does we can specify this at cost $O(1)$ but if not we can specify it at cost $O(\log n)$ as it is alive in $\rho^{*}$.
For each information piece in $J_{i}^{j}$ we thus have, in the worst case, to discover one mini-square at cost $t$, one at cost $\Delta$ and then six mini-squares at cost $O(\log n)$. This are two partners in $I_{i}^{j}$ and then up to four partners in $I^{*j+}$. This implies that the total external cost at stage $j$ is at most $2^{O(s_{j})}(\log n)^{6s_{j}}(t\Delta)^{s_{j}}$.
As any edge in $J_{i}^{j}$ might result in four queries in the common decision tree, if we have $4s$ questions in the this tree, then $\sum s_{j}\geq s$ and let us for convenience assume we have equality. The total cost of external information is in such a case $2^{O(s)}(\log n)^{6s}(t\Delta)^{s}$.
Lemma 4.9 applies also in this case from a similar calculation to that of the standard switching lemma we conclude that the total probability of quadruples $\tau$, $U$ $\pi_{2}$, $B$ that can reconstructed for a single choice of which formulas to process is bounded by
$M^{4s/\ell}\Delta 2^{O(s)}((\log n)^{7}t\Delta^{-1})^{s}.$
This completes the proof.
7 The lower bound for number of lines
We finally arrive at our second main theorem.
Theorem 7.1
Assume that $d\leq O(\frac{\log n}{\log{\log n}})$ and let $n$ be an odd integer and $M$ a parameter. Then any depth-$d$ Frege proof of the functional and onto PHP on the $n\times n$ grid where each line is of size $M$ requires at least
$\mbox{exp}\left(\Omega\left(\frac{n}{(\log M)^{2d-1}(\log n)^{O(1)}}\right)\right)$
lines.
Proof:  Given that we established the multi-switching lemma the situation is now essentially the same as in [11].
Parameters are similar as in the proof for the total proof size. We use $\ell=\log M$ in all applications of the Lemma 6.1 while we have $t=1$ in the first application and in later applications $t=\ell$. With these choices $M^{4s/\ell}$ equals $2^{O(s)}$ and can be included in the constant $A$.
We set $\Delta=\Theta((\log n)^{7}t)$ to ensure that the base of the exponent (including the contribution from the $M^{4s/\ell}$ factor) is bounded by $\frac{1}{32}$. Now suppose the proof has $N$ lines.
In our first application we set $4s_{1}=\log N$ and we conclude, by the union bound, that each line with high probability admits its own own $s_{1}$-common $\ell$-evaluation.
In the second iteration we have $2^{s_{1}}N$ sets of formulas to consider. This is the case as we need to consider each leaf of each decision tree in each line. By setting $s_{2}=2s_{1}$ we can apply the union bound and conclude that with high probability all depth 2 formulas are now in the range of the respective $\ell$-evaluations.
Continuing similarly, in the $i$th iteration we use $s_{i}=2^{i-1}\log N$. The final common decision tree is of of depth at most $\sum_{j=0}^{i-1}s_{j}=O(2^{d}\log N)$.
After $d$ rounds of restrictions we have reduced the size of the PHP from $n$ to $n/(\log n)^{O(d)}(\log M)^{2d-2}$. If this is larger than $150(\sum_{j=0}^{d}s_{j}+\log M)$ we can conclude that there is no proof and this gives the bound of the theorem.
8 Final words
It seems that to be able to prove a switching lemma for a space of restrictions one essential property is that for any given variable, the probability of setting it to either constant should be significantly higher than keeping it undetermined. In the unrestricted PHP the probability that any variable is true is about $\frac{1}{n}$ and to make the probability of the variables remaining undetermined smaller we must go from size $n$ to a size smaller than $\sqrt{n}$. With such a quick decrease in size we can only apply the switching lemma $O(\log\log n)$ times. In the graph PHP on the grid the probability of a variable being true to about $\frac{1}{4}$ and at the same time the probability of keeping a variable undetermined is about $\log n/\Delta$. It is harder to preserve a graph PHP, but this is made possible by using augmenting paths as the new variables. Thus it seems that both decision were forced upon us but certainly there might be other possibilities and it would be interesting to see alternative proofs for a switching lemma for a space of restrictions that preserve the PHP.
In this paper we have proved yet another switching lemma and it might not even be the last one. There are many remaining questions in both proof complexity and circuit complexity and some might be attackable by these types of techniques. It would be very interesting, however, to find a different technique to attack questions of Frege proofs where each formula is relatively simple.
While proofs of switching lemmas are non-trivial, the properties of the proof we use are rather limited. In the assumed proof of contradiction, after the restrictions, the proof, more or less, only contain formulas of constant size. It is not difficult to see that such a proof cannot find a contradiction for a set of axioms that are locally consistent. It would be interesting with a reasoning that used more interesting properties of being a proof.
Acknowledgments
I am deeply grateful to Svante Janson for suggesting the proof of Lemma 3.7. I am also very grateful to Per Austrin, Kilian Risse, Ben Rossman, and Aleksa Stanković for a sequence of discussions that lead to the start of the ideas for this paper. Kilian has also greatly helped to improve the quality of the writing of this paper. This is partly due to discussions about writing [11], but the also when discussing details of this paper. Finally, I thank the referees of this paper for several suggestions how to improve the presentation of the paper. The research leading to this publication was supported by the Knut and Alice Wallenberg foundation.